Determine the equation of the hyperbola with center (0,0), vertices (-5,0) and (5,0) and foci (-7,0)...

  1. Determine the equation of the hyperbola with center (0,0), vertices (-5,0) and (5,0) and foci (-7,0) and (7,0)
Answer #1

clue) Egn D 22 2) - ( il Q = 5 C= K-5,0) (5,0) (4,0) c² = a + b 2 y2 s²62 2 (-10) 6² =24 c Eqm x² -9-0 = 1 24 *